OpenStudy (anonymous):

sin^2(y) + 2sinycosy = 0. What do I do?? do i turn 2sinycosy to sin2y?? but how will that help?? :P

OpenStudy (anonymous):

@hartnn heeelp :P

hartnn (hartnn):

can't u take sin x common ?

hartnn (hartnn):

sin2y won't help

hartnn (hartnn):

sin y (sin y+2 cos y)=0

OpenStudy (phi):

factor out sin(y) then you have sin(y) ( sin(y) +2 cos(y)) =0 that means either sin(y)=0 or sin(y)+2cos(y)=0 this last one can be changed into a tan(y)

OpenStudy (anonymous):

yeahh, i thought of that :P but then what??

OpenStudy (anonymous):

HOLD UP how??

hartnn (hartnn):

@phi explained

OpenStudy (phi):

divide by cos(y)

hartnn (hartnn):

how for which part ?

OpenStudy (anonymous):

oh waitt, i think i got it :P

OpenStudy (anonymous):

lemme put in the answers in a few, tell me if its right :P same domain tho, btw.

OpenStudy (anonymous):

so i got 0, 180, 360, 116.57, 296.57?? :P thankss, guys :)

hartnn (hartnn):

yup, even i got that.

OpenStudy (anonymous):

AHA i feel so ishmaaart. shmarty-pantsh.

hartnn (hartnn):

lol! u are :)

OpenStudy (anonymous):

heyy, can you come to this question?? i need some help explaining this thing :P http://openstudy.com/study#/updates/50645c64e4b0583d5cd3c53c